OpenStudy (ayyookyndall):

Given triangle RST = triangle NPQ, RT = 7x - 5, NQ = 5x + 11, find the length of RT and NQ. a. 8 b. 51 c. 40 d. 56 ** I think it is B

OpenStudy (anonymous):

The answer is 8. You have to plug in the numbers and get the same answer. Plug in 8 and you'll get 7(8)=56-5=51 5(8)=40+11=51

OpenStudy (ayyookyndall):

Okay. Thank you!

OpenStudy (anonymous):

You're welcome. (: